OpenStudy (anonymous):

if : 2x=8^(x+1), value of x?

OpenStudy (anonymous):

Are you sure the left isn't 2^x?

OpenStudy (anonymous):

you get paid for this

OpenStudy (anonymous):

I'm employed by OpenStudy.

OpenStudy (anonymous):

o shoot ye its 2^x

OpenStudy (anonymous):

i jsut signed up today so i had no idea tatas nice

OpenStudy (anonymous):

2^x = 2^(3(x+1)) x=3x+3 2x=-3 x=-3/2

OpenStudy (anonymous):

k my last question if : 5^(3x*2)=25 what is x?

OpenStudy (anonymous):

5^(3x*2)=5^2 6x=2 x=1/3